# Under-coverage in high-statistics counting experiments with finite MC samples

### v2

@@ 2.3 Numerical results
1) Gauss (asympt.) Analytic Hessian 0 $0.683$ 0.094, $,\;0.094$ 0.094
$+5\bar{\sigma}_{{\rm H}}$ $0.683$ 0.094, $,\;0.094$ 0.094
$-5\bar{\sigma}_{{\rm H}}$ $0.683$ 0.094, $,\;0.094$ 0.094
2) Poisson Numeric Hessian 0 $0.516(5)$ 0.057, $,\;0.057$ 0.057
$+5\bar{\sigma}_{{\rm H}}$ $0.001(1)$ 0.057, $,\;0.057$ 0.057
$-5\bar{\sigma}_{{\rm H}}$ $0.002(1)$ 0.057, $,\;0.057$ 0.057
3) Gauss Numeric Hessian 0 $0.516(5)$ 0.057, $,\;0.057$ 0.057
$+5\bar{\sigma}_{{\rm H}}$ $0.001(1)$ 0.057, $,\;0.057$ 0.057
$-5\bar{\sigma}_{{\rm H}}$ $0.002(1)$ 0.057, $,\;0.057$ 0.057
4) Gauss Analytic Hessian 0 $0.516(5)$ 0.057, $,\;0.057$ 0.057
$+5\bar{\sigma}_{{\rm H}}$ $0.001(1)$ 0.057, $,\;0.057$ 0.057
$-5\bar{\sigma}_{{\rm H}}$ $0.002(1)$ 0.057, $,\;0.057$ 0.057
5) Gauss + MC stat. Analytic Hessian 0 ${\bf 0.675}(5)$ 0.080, $,\;0.080$ 0.080
$+5\bar{\sigma}_{{\rm H}}$ $0.004(1)$ 0.080, $,\;0.080$ 0.080
$-5\bar{\sigma}_{{\rm H}}$ $0.004(1)$ 0.080, $,\;0.080$ 0.080
6) Gauss + BB-lite Numeric Hessian 0 ${\bf 0.674}(5)$ 0.080, $,\;0.080$ 0.080
$+5\bar{\sigma}_{{\rm H}}$ $0.004(1)$ 0.080, $,\;0.080$ 0.080
$-5\bar{\sigma}_{{\rm H}}$ $0.004(1)$ 0.081, 0.0800.0810.0800.081,\;0.0800.081 , 0.080
7) Gauss + BB Numeric Hessian 0 $0.443(5)$ 0.139, 0.1330.1390.1330.139,\;0.1330.139 , 0.133
$+5\bar{\sigma}_{{\rm H}}$ $0.468(5)$ 0.147, 0.1400.1470.1400.147,\;0.1400.147 , 0.140
$-5\bar{\sigma}_{{\rm H}}$ $0.465(5)$ 0.147, 0.1410.1470.1410.147,\;0.1410.147 , 0.141
8) Gauss + BB Numeric PLR 0 $0.448(5)$ 0.140, 0.1340.1400.1340.140,\;0.1340.140 , 0.134
$+5\bar{\sigma}_{{\rm H}}$ $0.474(5)$ 0.152, 0.1440.1520.1440.152,\;0.1440.152 , 0.144
$-5\bar{\sigma}_{{\rm H}}$ $0.470(5)$ 0.143, 0.1370.1430.1370.143,\;0.1370.143 , 0.137
9) Gauss + BB Numeric sPFC 0 $0.474(5)$ 0.158, 0.1440.1580.1440.158,\;0.1440.158 , 0.144
$+5\bar{\sigma}_{{\rm H}}$ $0.597(5)$ 0.211, 0.2000.2110.2000.211,\;0.2000.211 , 0.200
$-5\bar{\sigma}_{{\rm H}}$ $0.590(5)$ 0.191, 0.1840.1910.1840.191,\;0.1840.191 , 0.184
10) Gauss + BB Numeric PFC 0 $0.471(5)$ 0.172, 0.1560.1720.1560.172,\;0.1560.172 , 0.156
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.671}(5)$ 0.222, 0.1960.2220.1960.222,\;0.1960.222 , 0.196
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.675}(5)$ 0.212, 0.1920.2120.1920.212,\;0.1920.212 , 0.192
11) Gauss + BB Numeric Barlett 0 $0.466(5)$ 0.167, 0.1540.1670.1540.167,\;0.1540.167 , 0.154
$+5\bar{\sigma}_{{\rm H}}$ $0.633(5)$ 0.213, 0.1880.2130.1880.213,\;0.1880.213 , 0.188
$-5\bar{\sigma}_{{\rm H}}$ $0.635(5)$ 0.201, 0.1810.2010.1810.201,\;0.1810.201 , 0.181
12) Gauss + BB Numeric CH 0 ${\bf 0.675}(5)$ 0.235, 0.2200.2350.2200.235,\;0.2200.235 , 0.220
$+5\bar{\sigma}_{{\rm H}}$ $0.884(3)$ 0.345, 0.3520.3450.3520.345,\;0.3520.345 , 0.352
$-5\bar{\sigma}_{{\rm H}}$ $0.884(3)$ 0.347, 0.3540.3470.3540.347,\;0.3540.347 , 0.354
13) Gauss + BB Numeric FC Cheat 0 ${\bf 0.677}(5)$ 0.228, 0.2060.2280.2060.228,\;0.2060.228 , 0.206
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.682}(5)$ 0.232, 0.2100.2320.2100.232,\;0.2100.232 , 0.210
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.672}(5)$ 0.221, 0.1990.2210.1990.221,\;0.1990.221 , 0.199
14) Gauss + BB Numeric Heuristic 0 ${\bf 0.675}(5)$ 0.229, 0.2220.2290.2220.229,\;0.2220.229 , 0.222
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.703}(5)$ 0.241, 0.2320.2410.2320.241,\;0.2320.241 , 0.232
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.700}(5)$ 0.242, 0.2330.2420.2330.242,\;0.2330.242 , 0.233
@@
Table 3: Coverage, mean, and median of the $1\sigma$ confidence interval for $\mu=\mu^{\prime}_{1}$ for the toy model with $N=2\times 10^{6}$, $n=200$, $k=1$, and $\epsilon=0.03$, obtained from an ensemble of $10^{4}$ identically repeated pseudo-experiments. Cases where the observed coverage agrees with the expectation of $68.3\%$ within a relative ±$\pm 5\%$ tolerance are highlighted with a bold font. When present, the number within parenthesis refers to the statistical error on the last digit.
@@
    Hessian ($7^{\rm th}$ row) or profile-likelihood ratio ($8^{\rm th}$ row) uncertainties are similar, showing that for each pseudo-data the log-likelihood function is reasonably parabolic around the minimum. However, the sampling distribution of $t_{\mu}$ is remarkably different from a chi-square with one degree of freedom, as illustrated by figure 2 for $\mu\equiv\mu_{1}^{\prime}=0$ on pseudo-data generated at the true value $\mu_{\rm t}=0$. The sampling distributions of the MLE estimator $\hat{\mu}$ and of the estimated Hessian uncertainty $\hat{\sigma}_{\rm H}$ are shown in figure 3. They both deviate visibly from a Gaussian distribution, with the latter featuring a long tail towards the right. Similar results are obtained when considering the other parameterization.
@@
Figure 2: Sampling distribution of the PLR function $t_{\mu}$ evaluated at $\mu\equiv\mu^{\prime}_{1}=0$ for data distributed according to the nominal model (i.e. with $\mu_{\rm t}=0$) with $N=2\times 10^{6}$, $\epsilon=0.03$, $n=200$, and $k=1$, obtained from an ensemble of $10^{4}$ identically repeated pseudo-experiments.
@@ 2.4 Beyond the asymptotic approximation
We have considered the following CI-setting methods inspired by the classical Feldman and Cousins (FC) unified approach FC . The main feature common to all methods is their use of the profile-likelihood ratio as an ordering principle with no priori assumption on its statistical distribution. The latter is indeed determined a posteriori from a collection of randomly generated pseudo-experiments. At variance with the original Feldman-Cousins formulation, the main challenge here is represented by the presence of nuisance parameters.
@@
    If the asymptotic approximation is not valid, the profile-likelihood ratio $t_{\mu}$ in eq. (16) is not necessarily chi-square distributed and the coverage properties of eq. (17) are not guaranteed. In the so-called profiled Feldman-Cousins approach acero2024 , a best guess of the nuisance parameters $\hat{\boldsymbol{\theta}}_{\mu}$ is first obtained by minimizing the NLL function at a fixed value of $\mu$. Then, for each point $\mu$ sampling the range of values that we wish to include/exclude from the confidence interval, an ensemble of so-called Feldman-Cousins pseudo-experiments FC are generated by drawing event counts $y_{i}$ around $f_{i}(\mu,\hat{\boldsymbol{\theta}}_{\mu})$. The sampling distribution $P_{\rm FC}$ of $t_{\mu}$ is then used to determine a critical value $c_{\mu}(\alpha)$ at significance level $\alpha$, such that
[…]
@@
    We considered an alternative construction of the FC interval where the best-fit estimators $(\hat{\mu},\hat{\boldsymbol{\theta}})$ are first determined for all parameters, including the POI; then, an ensemble of FC pseudo-experiments are generated by drawing Poisson distributed event counts around $f_{i}(\hat{\mu},\hat{\boldsymbol{\theta}})$; finally, the sampling distribution of $t_{\mu}$ and critical value $c_{\mu}$ are computed for $\mu=\hat{\mu}$. The resulting interval, which will be denoted as simplified FC, is then defined similarly to eq. (20) with $\mu_{k}=\hat{\mu}$.
@@
    In case the asymptotic approximation is not valid, it has been suggested by Barlett Barlett ; CowanUnc to modify the profile-likelihood ratio $t_{\mu}$ such that its expectation value equals its asymptotic value of one (for one degree of freedom), that is $t_{\mu}\to t^{\prime}_{\mu}=t_{\mu}/\langle t_{\mu}\rangle$. Thus, starting from the sampling distributions obtained from the FC pseudo-experiments, $\langle t_{\mu}\rangle$ is determined for $\mu\in\{1,1\pm 5\bar{\sigma}_{{\rm H}}\}$, and the corresponding confidence interval defined as
@@
    In the profiled FC approach, all nuisance parameters, including those related to the MC statistical uncertainties, are fixed to their best-fit to the data, but their post-fit uncertainty is neglected when the FC pseudo-experiments are generated. This can be partly accounted for by following an hybrid Bayesian-frequentist approach first proposed by Cousins and Highland COUSINS1992331 . In the Cousins-Highland (CH) approach, the bin contents $y_{i}$ in each FC pseudo-experiment are drawn from different central values $f_{i}(\mu,\tilde{\boldsymbol{\theta}}_{\mu})$, where $\tilde{\boldsymbol{\theta}}_{\mu}$ is sampled from a multivariate normal distribution with mean $\hat{\boldsymbol{\theta}}_{\mu}$ and covariance matrix given by the a posteriori covariance matrix of $\hat{\boldsymbol{\theta}}_{\mu}$. The resulting sampling distribution $P_{\rm HC}$ differs from $P_{\rm FC}$, generally leading to more conservative intervals.
@@
Using the numerical setup described in section 2.2, we studied the performances of the newly introduced CI-setting methods. All results are reported in tables 2 and 3. The main observations can be summarized as follows.
@@ 2.5 Recovering the correct coverage
    Increasing the size of the MC samples. Tables 5 and 6 collect the results obtained by increasing the statistical power of the MC simulation relative to the data ($k$) by factors of $10$ and $40$, respectively. As one would have naively expected, the impact of the limited MC statistics becomes increasingly less relevant as the MC sample size grows compared to the data. For $k=10$, the correct behavior is still not fully achieved, while for $k=40$ all likelihood functions and CI setting methods provide roughly the same intervals and the right coverage is recovered for all cases. This shows that the asymptotic properties of the MLE can be ultimately recovered once the MC template fluctuations become sufficiently small.
@@
    Improving the signal-to-background separation. The results reported in tables 7 and 8 have been obtained by choosing $\epsilon=0.06$ and $\epsilon=0.12$, respectively. As the value of $\epsilon$ increases, the model becomes better behaved, as also reflected by a reduction in the condition number of the Hessian matrix. The coverage of the Barlow-Beeston likelihood with Hessian or PLR uncertainties improves, but the correct coverage cannot be recovered even for $\epsilon\sim{\cal O}(10\%)$. The level of under-coverage from the FC methods remains stable.
@@
    Reducing the number of bins. The results reported in tables 9 and 10 have been obtained by rebinning the histograms by factors of $2$ and $10$, respectively. As $n$ gets smaller, the coverage of the Hessian- or PLR-method improves steadily, but the intervals slightly under-cover even for $n=20$. As expected, fluctuations of MC templates across bins tend to average out with a coarser binning. Again, the level of under-coverage from the FC method does not change significantly.
@@
    Increasing the size of both data and simulation. The results in tables 11 and 12 have been obtained by increasing the number of events in data by factors of $10$ and $100$, respectively, while keeping the same proportion between data and simulation. For sufficiently large values of $N$, the asymptotic limit is ultimately reached when using the full Barlow-Beeston likelihood function (7), so that CI-setting methods based on the Hessian uncertainty or the PLR can finally provide the right coverage. The FC-methods still suffer from systematic under-coverage.

### v3

@@ 2.5 Numerical results
1) Gauss (asympt.) Analytic Hessian 0 $0.683$ 0.094, 0.0940.094,\;0.094
$5\bar{\sigma}_{{\rm H}}$ $0.683$ 0.094, 0.0940.094,\;0.094
$5\bar{\sigma}_{{\rm H}}$ $0.683$ 0.094, 0.0940.094,\;0.094
2) Poisson Numeric Hessian 0 $0.516(5)$ 0.057, 0.0570.057,\;0.057
$5\bar{\sigma}_{{\rm H}}$ $0.001(1)$ 0.057, 0.0570.057,\;0.057
$5\bar{\sigma}_{{\rm H}}$ $0.002(1)$ 0.057, 0.0570.057,\;0.057
3) Gauss Numeric Hessian 0 $0.516(5)$ 0.057, 0.0570.057,\;0.057
$5\bar{\sigma}_{{\rm H}}$ $0.001(1)$ 0.057, 0.0570.057,\;0.057
$5\bar{\sigma}_{{\rm H}}$ $0.002(1)$ 0.057, 0.0570.057,\;0.057
4) Gauss Analytic Hessian 0 $0.516(5)$ 0.057, 0.0570.057,\;0.057
$5\bar{\sigma}_{{\rm H}}$ $0.001(1)$ 0.057, 0.0570.057,\;0.057
$5\bar{\sigma}_{{\rm H}}$ $0.002(1)$ 0.057, 0.0570.057,\;0.057
5) Gauss + MC stat. Analytic Hessian 0 ${\bf 0.675}(5)$ 0.080, 0.0800.080,\;0.080
$5\bar{\sigma}_{{\rm H}}$ $0.004(1)$ 0.080, 0.0800.080,\;0.080
$5\bar{\sigma}_{{\rm H}}$ $0.004(1)$ 0.080, 0.0800.080,\;0.080
6) Gauss + BB-lite Numeric Hessian 0 ${\bf 0.674}(5)$ 0.080, 0.0800.080,\;0.080
$5\bar{\sigma}_{{\rm H}}$ $0.004(1)$ 0.080, 0.0800.080,\;0.080
$5\bar{\sigma}_{{\rm H}}$ $0.004(1)$ 0.081, 0.0800.081,\;0.080
7) Gauss + BB Numeric Hessian 0 $0.443(5)$ 0.139, 0.1330.139,\;0.133
$5\bar{\sigma}_{{\rm H}}$ $0.468(5)$ 0.147, 0.1400.147,\;0.140
$5\bar{\sigma}_{{\rm H}}$ $0.465(5)$ 0.147, 0.1410.147,\;0.141
8) Gauss + BB Numeric PLR 0 $0.448(5)$ 0.140, 0.1340.140,\;0.134
$5\bar{\sigma}_{{\rm H}}$ $0.474(5)$ 0.152, 0.1440.152,\;0.144
$5\bar{\sigma}_{{\rm H}}$ $0.470(5)$ 0.143, 0.1370.143,\;0.137
9) Gauss + BB Numeric sPFC 0 $0.474(5)$ 0.158, 0.1440.158,\;0.144
$5\bar{\sigma}_{{\rm H}}$ $0.597(5)$ 0.211, 0.2000.211,\;0.200
$5\bar{\sigma}_{{\rm H}}$ $0.590(5)$ 0.191, 0.1840.191,\;0.184
10) Gauss + BB Numeric PFC 0 $0.471(5)$ 0.172, 0.1560.172,\;0.156
$5\bar{\sigma}_{{\rm H}}$ ${\bf 0.671}(5)$ 0.222, 0.1960.222,\;0.196
$5\bar{\sigma}_{{\rm H}}$ ${\bf 0.675}(5)$ 0.212, 0.1920.212,\;0.192
11) Gauss + BB Numeric Barlett 0 $0.466(5)$ 0.167, 0.1540.167,\;0.154
$5\bar{\sigma}_{{\rm H}}$ $0.633(5)$ 0.213, 0.1880.213,\;0.188
$5\bar{\sigma}_{{\rm H}}$ $0.635(5)$ 0.201, 0.1810.201,\;0.181
12) Gauss + BB Numeric CH 0 ${\bf 0.675}(5)$ 0.235, 0.2200.235,\;0.220
$5\bar{\sigma}_{{\rm H}}$ $0.884(3)$ 0.345, 0.3520.345,\;0.352
$5\bar{\sigma}_{{\rm H}}$ $0.884(3)$ 0.347, 0.3540.347,\;0.354
13) Gauss + BB Numeric FC Cheat 0 ${\bf 0.677}(5)$ 0.228, 0.2060.228,\;0.206
$5\bar{\sigma}_{{\rm H}}$ ${\bf 0.682}(5)$ 0.232, 0.2100.232,\;0.210
$5\bar{\sigma}_{{\rm H}}$ ${\bf 0.672}(5)$ 0.221, 0.1990.221,\;0.199
14) Gauss + BB Numeric Heuristic 0 ${\bf 0.675}(5)$ 0.229, 0.2220.229,\;0.222
$5\bar{\sigma}_{{\rm H}}$ ${\bf 0.703}(5)$ 0.241, 0.2320.241,\;0.232
$5\bar{\sigma}_{{\rm H}}$ ${\bf 0.700}(5)$ 0.242, 0.2330.242,\;0.233
@@
Table 4: Coverage, mean, and median of the $1\sigma$ confidence interval for $\mu=\mu^{\prime}_{1}$ for the toy model with $N=2\times 10^{6}$, $n=200$, $k=1$, and $\epsilon=0.03$, obtained from an ensemble of $10^{4}$ identically repeated pseudo-experiments. Cases where the observed coverage agrees with the expectation of $68.3\%$ within a relative $\pm 5\%$ tolerance are highlighted with a bold font. When present, the number within parenthesis refers to the statistical error on the last digit.
@@
    Uncertainties based on either the Hessian matrix ($7^{\rm th}$ row) or the profile-likelihood ratio ($8^{\rm th}$ row) in the Gaussian approximation are similar, showing that for each pseudo-data the log-likelihood function is reasonably parabolic around the minimum. However, the sampling distribution of $t_{\mu}$ is remarkably different from a chi-square with one degree of freedom, as illustrated by figure 2 for $\mu\equiv\mu_{1}^{\prime}=0$ on pseudo-data generated at the true value $\mu_{\rm t}=0$. The sampling distributions of the MLE estimator $\hat{\mu}$ and of the estimated Hessian uncertainty $\hat{\sigma}_{\rm H}$ are shown in figure 3. They both deviate visibly from a Gaussian distribution, with the latter featuring a long tail towards the right. Similar results are obtained when considering the other parameterization.
@@
Figure 2: Sampling distribution of $t_{\mu}$, the profile-likelihood function in the Gaussian approximation, evaluated at $\mu\equiv\mu^{\prime}_{1}=0$ for data distributed according to the nominal model (i.e. with $\mu^{\prime}_{\rm t}=0$) with $N=2\times 10^{6}$, $\epsilon=0.03$, $n=200$, and $k=1$, obtained from an ensemble of $10^{4}$ identically repeated pseudo-experiments.
@@ 2.6 Beyond the asymptotic approximation
We have considered the following CI-setting methods inspired by the classical Feldman and Cousins (FC) unified approach [15]. The main feature common to all methods is their use of the profile-likelihood ratio as an ordering principle with no priori assumption on its statistical distribution. The latter is indeed determined a posteriori from a collection of randomly generated pseudo-experiments. At variance with the original Feldman-Cousins formulation, the main challenge here is represented by the presence of nuisance parameters.
  • 1.
@@
    If the asymptotic approximation is not valid, the profile-likelihood ratio $t_{\mu}$ in eq. (19) is not necessarily chi-square distributed and the coverage properties of eq. (20) are not guaranteed. In the so-called profiled Feldman-Cousins approach [20], a best guess of the nuisance parameters $\hat{\boldsymbol{\theta}}_{\mu}$ is first obtained by minimizing the NLL function at a fixed value of $\mu$. Then, for each point $\mu$ sampling the range of values that we wish to include/exclude from the confidence interval, an ensemble of so-called Feldman-Cousins pseudo-experiments [15] are generated by drawing event counts $y_{i}$ around $f_{i}(\mu,\hat{\boldsymbol{\theta}}_{\mu})$. The sampling distribution $P_{\rm FC}$ of $t_{\mu}$ is then used to determine a critical value $c_{\mu}(\alpha)$ at significance level $\alpha$, such that
    $\int_{0}^{c_{\mu}(\alpha)}dt_{\mu}\,P_{\rm FC}(t_{\mu})=\alpha.$ (22)
@@
[…]
  • 3.
@@
    In case the asymptotic approximation is not valid, it has been suggested by Barlett [27, 11] to modify the profile-likelihood ratio $t_{\mu}$ such that its expectation value equals its asymptotic value of one (for one degree of freedom), that is $t_{\mu}\to t^{\prime}_{\mu}=t_{\mu}/\langle t_{\mu}\rangle$. Thus, starting from the sampling distributions obtained from the FC pseudo-experiments, $\langle t_{\mu}\rangle$ is determined for $\mu\in\{1,1\pm 5\bar{\sigma}_{{\rm H}}\}$, and the corresponding confidence interval defined as
@@
    In the profiled FC approach, all nuisance parameters, including those related to the MC statistical uncertainties, are fixed to their best-fit to the data, but their post-fit uncertainty is neglected when the FC pseudo-experiments are generated. This can be partly accounted for by following an hybrid Bayesian-frequentist approach first proposed by Cousins and Highland [8]. In the Cousins-Highland (CH) approach, the bin contents $y_{i}$ in each FC pseudo-experiment are drawn from different central values $f_{i}(\mu,\tilde{\boldsymbol{\theta}}_{\mu})$, where $\tilde{\boldsymbol{\theta}}_{\mu}$ is sampled from a multivariate normal distribution with mean $\hat{\boldsymbol{\theta}}_{\mu}$ and covariance matrix given by the a posteriori covariance matrix of $\hat{\boldsymbol{\theta}}_{\mu}$. The resulting sampling distribution $P_{\rm HC}$ differs from $P_{\rm FC}$, generally leading to more conservative intervals.
  • 5.
@@
Using the numerical setup described in section 2.4, we studied the performances of the newly introduced CI-setting methods. All results are reported in tables 3 and 4. The main observations can be summarized as follows.
@@ 2.7 Recovering the correct coverage
    Increasing the size of the MC samples. Tables 7 and 8 collect the results obtained by increasing the statistical power of the MC simulation relative to the data ($k$) by factors of $10$ and $40$, respectively. As one would have naively expected, the impact of the limited MC statistics becomes increasingly less relevant as the MC sample size grows compared to the data. For $k=10$, the correct behavior is still not fully achieved, while for $k=40$ all likelihood functions and CI setting methods provide roughly the same intervals and the right coverage is recovered for all cases. This shows that the asymptotic properties of the MLE can be ultimately recovered once the MC template fluctuations become sufficiently small.
@@
    Improving the signal-to-background separation. The results reported in tables 9 and 10 have been obtained by choosing $\epsilon=0.06$ and $\epsilon=0.12$, respectively. As the value of $\epsilon$ increases, the model becomes better behaved, as also reflected by a reduction in the condition number of the Hessian matrix. The coverage of the Barlow-Beeston likelihood with Hessian or PLR uncertainties improves, but the correct coverage cannot be recovered even for $\epsilon\sim{\cal O}(10\%)$. The level of under-coverage from the FC methods remains stable.
@@
    Reducing the number of bins. The results reported in tables 11 and 12 have been obtained by rebinning the histograms by factors of $2$ and $10$, respectively. As $n$ gets smaller, the coverage of the Hessian- or PLR-method improves steadily, but the intervals slightly under-cover even for $n=20$. As expected, fluctuations of MC templates across bins tend to average out with a coarser binning. Again, the level of under-coverage from the FC method does not change significantly.
@@
    Increasing the size of both data and simulation. The results in tables 13 and 14 have been obtained by increasing the number of events in data by factors of $10$ and $100$, respectively, while keeping the same proportion between data and simulation. For sufficiently large values of $N$, the asymptotic limit is ultimately reached when using the full Barlow-Beeston likelihood function (8), so that CI-setting methods based on the Hessian uncertainty or the PLR can finally provide the right coverage. The FC-methods still suffer from systematic under-coverage.